OpenStudy (anonymous):

Use the product rule to find derivative y=(7x-6)^2

OpenStudy (jdoe0001):

\(\bf y=(7x-6)^2\implies y=(7x-6)(7x-6)\\ \quad \\ \cfrac{d}{dx}[(7x-6)]\cdot (7x-6)+(7x-6)\cdot \cfrac{d}{dx}[(7x-6)]\)

OpenStudy (anonymous):

i got 182x^2 as answer

OpenStudy (zzr0ck3r):

why in gods name would you do that.

OpenStudy (jdoe0001):

heheh

OpenStudy (anonymous):

after that step that you did..i did (7x-0)(7x-0)

OpenStudy (jdoe0001):

\(\bf \cfrac{d}{dx}[7x]\implies 7\cdot 1x^{1-1}\)

OpenStudy (anonymous):

wait is it 98x^2+84x

OpenStudy (jdoe0001):

ahemm.... close

OpenStudy (jdoe0001):

\(\bf \cfrac{d}{dx}[7x]\implies 7\cdot 1x^{1-1}\implies 7x^0\implies 7\)

OpenStudy (anonymous):

98x^2+84

OpenStudy (jdoe0001):

98\(\large\bf x\)+84

OpenStudy (jdoe0001):

well. dohhh ahemm -84 rather

OpenStudy (jdoe0001):

check your distribution

OpenStudy (anonymous):

wait so whats the answer

OpenStudy (jdoe0001):

well... what's the derivative of (7x -6) ?

OpenStudy (anonymous):

y=49x^2+42+49x^2+42

OpenStudy (jdoe0001):

\(\bf y=(7x-6)^2\implies y=(7x-6)(7x-6)\\ \quad \\ {\color{blue}{ \cfrac{d}{dx}[(7x-6)]}}\cdot (7x-6)+(7x-6)\cdot {\color{blue}{ \cfrac{d}{dx}[(7x-6)]}}\\ \quad \\ {\color{green}{ \cfrac{d}{dx}[7x-6]\implies 7\cdot 1x^{1-1}-0\implies 7x^0\implies 7}}\qquad thus\\ \quad \\ {\color{blue}{ 7}}\cdot (7x-6)+(7x-6)\cdot {\color{blue}{7}}\)

OpenStudy (anonymous):

\[y=98x^2-0\]

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

hmm recheck your distribution

OpenStudy (anonymous):

omg i dont know what im doing wrong

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ 7}}\cdot (7x-6){\huge +}(7x-6)\cdot {\color{blue}{7}}\)

OpenStudy (anonymous):

49x-42+49x-42

OpenStudy (jdoe0001):

yeap